There is not much to report from the GreenHab today. We have moved our experiments into the lower level of the Hab and will not be able to move them back into the GreenHab until the heater is functioning and we have control of the temperature. Ideally, we would like to be able to maintain the temperature between 70 and 80 degrees F at all times. In contrast, the low temperature last night was 26 F and peaked today at 93 F. Obviously, these are not conditions that are suitable for plant growth. The first step in once again regaining control of the temperature is to get the heater functioning again.
The germination front is going quite well! We saw many more seedlings sprout and now have confirmation that all types of species that we planted at the beginning of our mission have germinated. We are very excited to see the pinto bean, kidney bean, and popcorn grow up before the end of our mission. Over the next few days, we will be working on transplanting the more mature seedlings and providing them nutrients through hydroponics and soil as well as beginning data collection on the lettuce experiments.
SEDS-MDRS Crew GreenHab Officer #2
SEDS-MDRS Crew HSO